A movie stunt driver on a motorcycle speeds horizontally off a 44.0 m high cliff. Determine the time it takes the motorcycle to leave the cliff top to land on the ground. Ignore air resistance. (g = 9.80 m/s2) Please explain.
Answer #1

as the speed of motorcycle is in horizontal direction, there is no vertical component of speed .

hence, only acceleration due to gravity is present in vertical direction.

applying 2nd equaion of motion,

s = u*t + 1/2*a*t2

where, u = 0 (in vertical direction)

s= 44 m

a = g= 9.8m/s2

so,

44 = 0+ 1/2*9.8*t2

t2 = 44*2/9.8 =8.979

t= 2.996 seconds